OpenStudy (yumeybitaa):

Evaluate [(21 + 6) - 32] ÷ 9 • 2. 4 4.6 1 52

OpenStudy (shadowlegendx):

NO ASTRO!!! mineEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEEE

OpenStudy (shadowlegendx):

BEDMAS, you know it? c;

OpenStudy (shadowlegendx):

BEDMAS order of operations. Each letter stands for: B - Brackets Exponents Division Multiplication Addition Subtraction

OpenStudy (shadowlegendx):

NUUH

OpenStudy (yumeybitaa):

can you please help me with my question or not ?

OpenStudy (shadowlegendx):

BEDMAS order of operations. Each letter stands for: B - Brackets Exponents Division Multiplication Addition Subtraction

OpenStudy (shadowlegendx):

You do the problem in Brackets first

OpenStudy (shadowlegendx):

So for this problem: [(21 + 6) - 32] ÷ 9 • 2. We would solve for: [(21 + 6) - 32] first

OpenStudy (shadowlegendx):

Since 21 + 6 is in parentheses in the brackets, we solve for that first

OpenStudy (shadowlegendx):

So what is 21 + 6?

OpenStudy (yumeybitaa):

27

OpenStudy (shadowlegendx):

Now we have: [(27) - 32] /9 x 2

OpenStudy (shadowlegendx):

/ = divide

OpenStudy (shadowlegendx):

So now we solve for 27 - 32, what is 27 - 32?

OpenStudy (astrophysics):

Wait lets clarify the question first \[\frac{ [(21+6)-32] }{ 9*2 }\] or \[\frac{ [(21+6)-32] }{ 9 }\times2\]

OpenStudy (yumeybitaa):

-5

OpenStudy (shadowlegendx):

@Astrophysics I'm not sure she will understand that concept

OpenStudy (astrophysics):

Lol alright, carry on XD

OpenStudy (shadowlegendx):

Alright, now we have: \[-5 \div 9 \times 2 \]

OpenStudy (shadowlegendx):

We solve for \[-5 \div 9 \]

OpenStudy (yumeybitaa):

0.55555556

OpenStudy (yumeybitaa):

??

OpenStudy (shadowlegendx):

@Yumeybitaa can you check and see if there are supposed to be any negative signs on the answers

OpenStudy (yumeybitaa):

oh yes forgot to put it

OpenStudy (shadowlegendx):

Alright so what is \[-0.5 \times 2 \]

OpenStudy (yumeybitaa):

1

OpenStudy (shadowlegendx):

A negative times a positive equals a negative

OpenStudy (shadowlegendx):

\[-0.5 \times 2 = -x\]

OpenStudy (shadowlegendx):

@Yumeybitaa -0.5 x 2?

OpenStudy (yumeybitaa):

-1 forgot to put the sign again

OpenStudy (shadowlegendx):

Yes, -1 is correct

OpenStudy (shadowlegendx):

This is a simple process of Order of Operations. Remember Brackets first, then Parentheses. You can use BEDMAS or PEMDAS. B = Brackets P = Parentheses D = Division M = Multiplication A = Addition S = Subtraction

OpenStudy (yumeybitaa):

thank you , i know that but in spanish

OpenStudy (shadowlegendx):

Oh ok :)